5. Solve the initial value problem, y+y - by = 4e* with y(0) = 1 and y(0) = 1
Answer #1

The find Yc Criven that, y+y- by = 40 40) , YO general solution is y=Yc+ Yo where Y c = complementary solution and Up = parUsing, 460) = 1 y(0)=1 y=-3c, é -3x 27 +2 e -e CI+(2-1 0 :. y 10) = 1 y lo-1 -36, + 2C2-1 = 1 2 Multiplying eq (1) by 3 an